# Exploratory study on the masses of odd-$Z$ nuclei and $r$-process simulation based on the deformed relativistic Hartree-Bogoliubov theory in continuum

### v1

@@ Abstract
Background: Nuclear masses of exotic nuclei are important for both nuclear physics and astrophysics. The deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) is capable of providing proper descriptions for exotic nuclei by simultaneously including deformation and continuum effects. The mass table of even-$Z$ nuclei with $8\leqslant Z\leqslant 120$ has been established based on the DRHBc theory[At. Data Nucl. Data Tables 158, 101661 (2024)].
Purpose: This work aims to systematically estimate the masses of odd-$Z$ nuclei based on the available DRHBc results of even-$Z$ nuclei, thereby providing a pseudo DRHBc mass table for all nuclei with $8\leqslant Z\leqslant 120$. This mass table can then be employed in the $r$-process studies to investigate the influence of deformation on $r$-process.
Method: The mass of an odd nucleus is expressed as a function of the masses and odd-even mass differences of its neighboring even nuclei, with the odd-even mass difference approximated by the average pairing gap. The $r$-process simulations are carried out using the site-independent classical $r$-process model based on the waiting-point approximation.
Results and Conclusions: The approximation of the odd-even mass difference with the average pairing gap is validated to be effective, by reproducing the masses of even-$Z$ odd-$N$ nuclei calculated by DRHBc. Combining the DRHBc masses of even-$Z$ nuclei and the estimated masses of odd-$Z$, a pseudo DRHBc mass table is established, with the root-mean-square (rms) deviation from available mass data $\sigma=1.50$ MeV. This pseudo DRHBc mass table is applied to the $r$-process simulation, and the impact of nuclear deformation effects is analyzed. The deformation effects can influence the $r$-process path and thus affect the $r$-process abundance. In particular, the nuclear shape transitions can even lead to the discontinuity of the $r$-process path.
I Introduction
@@
Nuclear masses are important for both nuclear physics [1] and astrophysics [2]. They reflect a variety of underlying physical effects of nuclear quantum many-body systems, and can be used to extract nuclear structure information, e.g., nuclear deformation [3], shell effects [4], and nuclear force [5]. They also determine the reaction energies for all nuclear reactions, which are important in understanding the energy production in stars [6] and the study of nucleosynthesis [7]. Experimentally, the masses of about 2500 nuclei have been measured so far [8].
@@
The rapid neutron capture process ($r$-process) has been believed to be responsible for the nucleosynthesis of about half of the elements heavier than iron for more than half a century [9, 10, 11]. The understanding of $r$-process is still affected by the uncertainties in our knowledge of both nuclear physics quantities and astrophysical conditions [12, 2, 11, 13, 14]. The masses of neutron-rich nuclei are crucial to the $r$-process studies, as they are needed to extract the reaction energies that go into the calculations of all involved nuclear reaction rates, i.e., neutron separation energies ($S_{n}$) for neutron-capture cross sections and $\beta$-decay $Q$-values ($Q_{\beta}$) for $\beta$-decay half-lives ($T_{1/2}$[15, 16]. Even in the classical $r$-process model with waiting-point approximation (WPA), the use of different nuclear mass models would affect the equilibrium between neutron capture and photodisintegration reactions to predict $r$-process path and also the $Q$-values of $\beta$-decay half-lives, and thus affect the final calculated abundance [17]. Nuclear masses of most neutron-rich nuclei, especially the exotic nuclei near drip lines, remain beyond the current experimental capabilities even in the foreseeable future, due to the difficulties in production, separation, and detection. Therefore, theoretical models with reliable predictive power are essential.
@@
Many efforts have been made to theoretically describe nuclear masses, including macroscopic-microscopic models [18, 19, 20, 21], microscopic models [22, 23, 24, 25, 26, 27], and machine-learning approaches [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Microscopic models are usually believed to have a better reliability of extrapolation [39, 40, 41]. This is particularly necessary for the mass predictions for the exotic nuclei, which are far away from the experimentally known region and play important roles in the $r$-process path.
@@
The exotic nuclei involved in the $r$-process path are extremely weakly bound, and their Fermi energies are very close to the continuum threshold. For these nuclei, the pairing interaction can scatter nucleons from bound states to the resonant ones in the continuum, and the density could become more diffuse due to this coupling to the continuum [42]. The stability of exotic nuclei and even the position of the drip line might be influenced, which is the so-called continuum effect. Therefore, the effects of pairing correlation and the coupling to continuum should be considered properly [43, 44, 45, 46, 47, 48] in the description for exotic nuclei near drip lines. This is important for the studies of $r$-process in extremely high-neutron-density environments, e.g., the neutron star mergers. The relativistic continuum Hartree-Bogoliubov (RCHB) theory [49, 50] takes into account pairing correlations and continuum effect in a microscopic and self-consistent way, and has achieved great success in describing both stable and exotic nuclei [51, 42]. Based on the RCHB theory, the first nuclear mass table including continuum effect was constructed [25], and the continuum effect on the limits of the nuclear landscape was studied. However, the accuracy of the RCHB mass table is limited due to the assumed spherical symmetry.
@@
Most nuclei in the nuclear chart, except for doubly magic ones, exhibit deformed shapes. In order to properly describe deformed exotic nuclei, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed by treating the effects of deformation, pairing correlations, and continuum simultaneously [52, 53]. As the advantages of the RCHB theory are inherited and the deformation degree of freedom is further included, the DRHBc theory has been successfully applied in a variety of studies on exotic nuclei, including the halo structures [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], peninsulas beyond the neutron drip line [40, 66, 67, 41], evolution of deformation and shape coexistence [68, 69, 70, 71], half-life estimation for proton emission and $\alpha$-decay [72, 73, 74], etc. The predictive power of the DRHBc theory for nuclear mass has been examined in Refs. [40, 41] by taking the even-even superheavy nuclei with $102\leqslant Z\leqslant 120$ as examples. Recently, based on the DRHBc theory, a nuclear mass table for the nuclei with $8\leqslant Z\leqslant 120$ is in progress [75, 76]. The DRHBc mass table is expected to provide the first microscopic mass input with the effects of deformation and continuum simultaneously included. Up to now, the even-$Z$ part of the DRHBc mass table has been established [27, 77].
@@
Since the results for odd-$Z$ nuclei are unavailable in the DRHBc mass table, our work aims to systematically estimate their masses based on the available results of even-$Z$ nuclei, and then utilize them to perform $r$-process simulations. This article is structured as follows: The theoretical frameworks for the DRHBc theory and classical $r$-process are briefly introduced in Section II, the masses of odd-$Z$ nuclei are estimated in Section III, the $r$-process simulation based on the obtained masses is discussed in Section IV, and finally, a summary is given in Section V.
II Theoretical framework
II.1 The DRHBc theory
@@
The details of the DRHBc theory have been illustrated in Refs. [53, 75, 76]. In this Section, we just present a brief theoretical framework. In the DRHBc theory, the motion of nucleons are microscopically described by the relativistic Hartree-Bogoliubov (RHB) equation [78],
@@
where $V^{pp}$ is the pairing force, and $\kappa$ is the pairing tensor [79]. In this work, the density-dependent zero-range pairing force
$V^{pp}(\bm{1}_{1},\bm{r}_{2})=V_{0}\frac{1}{2}(1-P^{\sigma})\delta(\bm{r}_{1}-\bm{r}_{2})\left(1-\frac{\rho(\bm{r}_{1})}{\rho_{\text{sat}}}\right)$ (4)
@@
$f(\bm{r})=\sum_{\lambda}f_{\lambda}(r)P_{\lambda}(\cos\theta),\qquad\lambda=0,2,4,\dots,\lambda_{\max}.$ (5)
For the exotic nuclei close to drip lines, the continuum effect should be taken into account properly [44, 42]. For this purpose, the deformed RHB equations (1) are solved in a spherical Dirac Woods-Saxon basis [80, 81], which can properly describe the asymptotic behavior of the density distribution at a large $r$ for exotic nuclei.
For a nucleus with odd number of neutron or proton, the blocking effect of the unpaired nucleon(s) needs to be considered [79]. Practically, this can be realized by the exchange of quasiparticle wavefunctions $(U_{k_{b}},V_{k_{b}})\leftrightarrow(V_{k_{b}}^{*},U_{k_{b}}^{*})$ and that of the energy $E_{k_{b}}\leftrightarrow-E_{k_{b}}$ for Eq. (1), where $k_{b}$ refers to the blocked orbital for the odd nucleon [82, 76].
After self-consistently solving the RHB equations, the expectation values such as binding energy, quadrupole deformation, root-mean-square radii, etc., can be calculated [53, 75, 76]. The canonical basis |$\ket{\psi_{i}}$ is obtained by the following diagonalization [79]:
$\hat{\rho}\ket{\psi_{i}}=v_{i}^{2}\ket{\psi_{i}},$ (6)
where $\hat{\rho}$ is the density matrix, and $v_{i}^{2}$ is the corresponding occupation probability of |$\ket{\psi_{i}}$. The single-particle energy in the canonical basis is obtained as $\epsilon_{i}=\braket{\psi_{i}}{\hat{h}_{D}}{\psi_{i}}$. The pairing gap $\Delta_{i}$ is calculated by
@@
where the parameter $u_{i}$ is obtained from $u_{i}^{2}+v_{i}^{2}=1$. The average pairing gap defined by [44, 83]:
@@
is an order parameter describing the phase transition from a normal fluid to a superfluid [84].
II.2 Classical $r$-process
In this work, a site-independent $r$-process model, i.e., the so-called classical $r$-process model, is employed to specifically study the effect of nuclear mass on the $r$-process simulation. Classical $r$-process model can be regarded as a realistic simplification of the dynamical $r$-process model, and has been successfully employed in describing $r$-process patterns of both the solar system and metal-poor stars [85, 86, 87, 88]. Nevertheless, it should be noted that the real neutron freeze-out after the equilibrium between neutron capture and photodisintegration reactions, as well as the fission recycling, are neglected in the present classical $r$-process model.
In the classical $r$-process model, iron group seed nuclei are irradiated by high-density neutron sources with a high temperature $T\gtrsim 1.5~{}\mathrm{GK}$. The $r$-process abundances are obtained by the superposition of abundances from the simulations in 16 different neutron flows with neutron densities in the range of $10^{20}-10^{27.5}\mathrm{~{}cm}^{-3}$. The weight $\omega$ and the irradiation time $\tau$ of each neutron flow follow exponential relations on neutron density $n_{n}$ [85, 89]:

### v2

@@ Abstract
Nuclear masses of exotic nuclei are important for both nuclear physics and astrophysics. The deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) is capable of providing proper descriptions for exotic nuclei by simultaneously including deformation, pairing correlation and continuum effects, and a mass table of even-$Z$ nuclei with $8\leqslant Z\leqslant 120$ has been developed based on the DRHBc theory. This work employs a methodology to estimate the masses of odd nuclei using neighboring even nuclei’s masses and microscopic pairing gaps, and the performance of microscopic pairing gaps are validated by comparing with empirical ones. Combining the DRHBc masses of even-$Z$ nuclei and the estimated masses of odd-$Z$ nuclei, a pseudo DRHBc mass table is developed, with the root-mean-square (rms) deviation from available mass data $\sigma=1.47$ MeV. Then this mass table is employed in the $r$-process simulation; results show that the differences in the details of pairing gaps do not yield qualitative discrepancy in $r$-process abundances, while the deformation effects can influence the $r$-process path and thus affect the $r$-process abundance. In particular, the nuclear shape transitions can even lead to the discontinuity of the $r$-process path, suggesting that incorporating triaxiality or beyond-mean-field effects would be valuable for further improvement.
1 Introduction
@@
Nuclear masses are important for both nuclear physics Yamaguchi2021PPNP and astrophysics Kajino2019PPNP . They reflect a variety of underlying physical effects of nuclear quantum many-body systems, and can be used to extract nuclear structure information, e.g., nuclear deformation Roubin2017PRC , shell effects Ramirez2012Science , and nuclear force Wienholtz2013Nature . They also determine the reaction energies for all nuclear reactions, which are important in understanding the energy production in stars Bethe1939PR and the study of nucleosynthesis Mumpower2015PRC . However, experimentally, the masses of about 2500 nuclei have been measured so far Wang2021CPC .
@@
The rapid neutron capture process ($r$-process) has been believed to be responsible for the nucleosynthesis of about half of the elements heavier than iron for more than half a century Burbidge1957RMP ; Thielemann2017ARNPS ; Cowan2021RMP . The understanding of $r$-process is still affected by the uncertainties in our knowledge of both nuclear physics quantities and astrophysical conditions Zhao2019APJ ; Kajino2019PPNP ; Cowan2021RMP ; Arcones2023AAR ; Wu2023Sci.Bull. . The masses of neutron-rich nuclei are crucial to the $r$-process studies, as they are needed to extract the reaction energies that go into the calculations of all involved nuclear reaction rates, i.e., neutron separation energies ($S_{n}$) for neutron-capture cross sections and $\beta$-decay $Q$-values ($Q_{\beta}$) for $\beta$-decay half-lives ($T_{1/2}$Mumpower2016PPNP ; Wu2022APJ ; Huang2025ApJ . Even in the classical $r$-process model with waiting-point approximation (WPA), the use of different nuclear mass models would affect the equilibrium between neutron capture and photodisintegration reactions to predict $r$-process path and also the $Q$-values of $\beta$-decay half-lives, and thus affect the final calculated abundance Jiang2021APJ . Nuclear masses of most neutron-rich nuclei, especially the exotic nuclei near drip lines, remain beyond the current experimental capabilities even in the foreseeable future, due to the difficulties in production, separation, and detection. Therefore, theoretical models with reliable predictive power are essential.
@@
Many efforts have been made to theoretically describe nuclear masses, including macroscopic-microscopic models Pearson1996PLB ; Wang2014PLB ; Moller2016ADNDT ; Koura2005PTP , microscopic models Goriely2009PRL_Skyrme ; Goriely2009PRL_Gogny ; Pena-Arteaga2016EPJA ; Xia2018ADNDT ; Yang2021PRC ; Zhang2022ADNDT , and machine-learning approaches Utama2016PRC ; Neufcourt2018PRC ; Neufcourt2019PRL ; Wu2020PRC ; Wu2021PLB ; Niu2022PRC ; Du2023CPC ; Wu2024PRC ; Wu2024PRC_AKRR ; Wu2024SC ; Guo2024PRC . Microscopic models are usually believed to have a better reliability of extrapolation Zhao2012PRCmass ; Zhang2021PRC ; He2024PRC . This is particularly necessary for the mass predictions for the exotic nuclei, which are far away from the experimentally known region and play important roles in the $r$-process path.
@@
The exotic nuclei involved in the $r$-process path are extremely weakly bound, and their Fermi energies are very close to the continuum threshold. For these nuclei, the pairing interaction can scatter nucleons from bound states to the resonant ones in the continuum, and the density could become more diffuse due to this coupling to the continuum Meng2006PPNP . The stability of exotic nuclei and even the position of the drip line might be influenced, which is the so-called continuum effect. Therefore, the effects of pairing correlation and the coupling to continuum should be considered properly Dobaczewski1984NPA ; Dobaczewski1996PRC ; Grasso2001PRC ; Michel2008PRC ; Pei2011PRC ; Zhang2013PRC in the description for exotic nuclei near drip lines. This is important for the studies of $r$-process in extremely high-neutron-density environments, e.g., the neutron star mergers. The relativistic continuum Hartree-Bogoliubov (RCHB) theory Meng1996PRL ; Meng1998NPA , by assuming spherical symmetry and solving the relativistic Hartree-Bogoliubov (RHB) equations for nucleons in the coordinate space, takes into account pairing correlations and continuum effect in a microscopic and self-consistent way, and has achieved great success in describing both stable and exotic nuclei Meng1998PRL ; Meng2006PPNP .
Based on the RCHB theory, the first nuclear mass table including continuum effect was constructed Xia2018ADNDT , and the continuum effect on the limits of the nuclear landscape was studied. However, the accuracy of the RCHB mass table is limited due to the assumed spherical symmetry.
@@
Most nuclei in the nuclear chart, except for those with shell closures, exhibit deformed shapes. In order to properly describe deformed exotic nuclei, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed by solving the deformed RHB equations in a Dirac Woods-Saxon basis, and can treat the effects of deformation, pairing correlations, and continuum simultaneously Zhou2010PRC ; Li2012PRC . As the advantages of the RCHB theory are inherited and the deformation degree of freedom is further included, the DRHBc theory has been successfully applied in a variety of studies on exotic nuclei, including the halo structures Sun2018PLB ; Yang2021PRL ; Sun2021PRC ; Zhong2022SCP ; Zhang2023PRC_Na ; Zhang2023PLB ; Pan2024PLB ; Zhang2024PRC_Al , peninsulas beyond the neutron drip line Zhang2021PRC ; Pan2021PRC ; He2021CPC ; He2024PRC , evolution of nuclear shape and size Choi2022PRC ; Kim2022PRC ; Mun2023PLB ; Guo2023PRC ; Pan2025PRC , half-life estimation for proton emission and $\alpha$-decay Xiao2023PLB ; Choi2024PRC ; Lu2024PLB , etc. The predictive power of the DRHBc theory for nuclear mass has been examined in Refs. Zhang2021PRC ; He2024PRC by taking the even-even superheavy nuclei with $102\leqslant Z\leqslant 120$ as examples. Recently, based on the DRHBc theory with the relativistic density functional PC-PK1 and the density-dependent zero-range pairing force, a nuclear mass table for the nuclei with $8\leqslant Z\leqslant 120$ is in progress Zhang2020PRC ; Pan2022PRC . The DRHBc mass table is expected to provide a microscopic mass input with the effects of deformation and continuum simultaneously included. Up to now, the even-$Z$ part of the DRHBc mass table has been developed Zhang2022ADNDT ; Guo2024ADNDT .
@@
Since the results for odd-$Z$ nuclei are unavailable in the current DRHBc mass table, our work aims to systematically estimate their masses based on the available results of even-$Z$ nuclei, and then utilize them to perform $r$-process simulations. This article is structured as follows: The theoretical frameworks for the DRHBc theory and classical $r$-process are briefly introduced in Section 2, the examination of the methodology and the estimation of the masses of odd-$Z$ nuclei are presented in Section 3, the $r$-process simulation based on the obtained masses and the impact of the deformation effects are discussed in Section 4, and finally, a summary is given in Section 5.
2 Theoretical framework
2.1 The DRHBc theory
@@
The details of the DRHBc theory have been illustrated in Refs. Li2012PRC ; Zhang2020PRC ; Pan2022PRC . In this Section, we just present a brief theoretical framework. In the DRHBc theory, the motion of nucleons are microscopically described by the relativistic Hartree-Bogoliubov (RHB) equation Kucharek1991ZPA ,
@@
where $V^{pp}$ is the pairing force, and $\kappa$ is the pairing tensor Ring1980NMBP . In this work, the density-dependent zero-range pairing force
$V^{pp}(\bm{r}_{1},\bm{r}_{2})=V_{0}\frac{1}{2}(1-P^{\sigma})\delta(\bm{r}_{1}-\bm{r}_{2})\left(1-\frac{\rho(\bm{r}_{1})}{\rho_{\text{sat}}}\right)$ (4)
@@
$f(\bm{r})=\sum_{\lambda}f_{\lambda}(r)P_{\lambda}(\cos\theta),\quad\lambda=0,2,4,\dots,\lambda_{\max}.$ (5)
For the exotic nuclei close to drip lines, the continuum effect should be taken into account properly Dobaczewski1996PRC ; Meng2006PPNP . For this purpose, the deformed RHB equations (1) are solved in a spherical Dirac Woods-Saxon basis Zhou2003PRC ; Zhang2022PRC , which can properly describe the asymptotic behavior of the density distribution at a large $r$ for exotic nuclei. Since the odd nucleon breaks the time-reversal symmetry and leads to nonzero nucleon currents Pan2024PLB , to avoid such complexity, the time-reversal symmetry is assumed with the equal filling approximation Perez-Martin2008PRC ; Li2012CPL .
For a nucleus with odd number of neutron or proton, the blocking effect of the unpaired nucleon(s) needs to be considered Ring1980NMBP . Practically, this can be realized by the exchange of quasiparticle wavefunctions $(U_{k_{b}},V_{k_{b}})\leftrightarrow(V_{k_{b}}^{*},U_{k_{b}}^{*})$ and that of the energy $E_{k_{b}}\leftrightarrow-E_{k_{b}}$ for Eq. (1), where $k_{b}$ refers to the blocked orbital for the odd nucleon Li2012CPL ; Pan2022PRC .
[…]
@@
where the parameter $u_{i}$ is obtained from $u_{i}^{2}+v_{i}^{2}=1$. The average pairing gap defined by Dobaczewski1996PRC ; Agbemava2014PRC :
@@
is an order parameter describing the phase transition from a normal fluid to a superfluid Afanasjev2015PRC . It should also be mentioned that for deformed nuclei, due to the breaking of the rotational symmetry, one should also consider the rotational correction energy Zhao2010PRC ; Zhang2020PRC ,
$E_{\mathrm{rot}}=-\frac{\braket{\hat{\bm{J}}^{2}}}{2\mathscr{I}},$ (9)
where $\hat{\bm{J}}$ is the total angular momentum operator, and $\mathscr{I}$ is the moment of inertia calculated by the Inglis-Belyaev formula with the cranking approximation Ring1980NMBP . In Refs. Zhang2022ADNDT ; Guo2024ADNDT , it was mentioned that the cranking approximation is not suitable for (near-)spherical nuclei, and therefore, the $E_{\text{rot}}$ values of the nuclei with $|\beta_{2}|<0.05$ are taken as zero in the DRHBc mass table. The two-dimensional collective Hamiltonian method Sun2022CPC could better consider the correction energies for the near-spherical nuclei Zhang2023PRC_2DCH and is expected to provide an alternative treatment for future improvement for the DRHBc mass table.
For completeness, here we briefly list the numerical conditions adopted in the DRHBc mass table calculations Zhang2020PRC ; Pan2022PRC : For the particle-hole channel, the relativistic density functional PC-PK1 Zhao2010PRC is employed, which can provide one of the best density-functional descriptions for nuclear masses Zhao2012PRCmass ; Zhang2014FoP ; Lu2015PRC ; Zhang2021PRC . For the particle-particle channel, a density-dependent zero-range pairing force (4) is adopted, with the pairing strength $V_{0}=-325~\mathrm{MeV~fm}^{3}$, the saturation density of nuclear matter $\rho_{\text{sat}}=0.152~\mathrm{fm}^{-3}$, and a pairing window 100 MeV, which are fixed by reproducing the odd-even mass difference data for calcium and lead isotopes Zhang2020PRC . The box size is $R_{\text{box}}=20$ fm and the mesh size is $\Delta r=0.1$ fm. For the Dirac Woods-Saxon basis, the angular momentum cutoff is $J_{\max}=23/2$$\hbar$, the energy cutoff is $E_{\text{cut}}^{+}=300$ MeV, and the number of the basis states in the Dirac sea is the same as that in the Fermi sea Zhou2003PRC . The Legendre expansion truncation for potentials and densities in Eq. (5) is $\lambda_{\max}=6$, 8 and 10 for the nuclei with $8\leqslant Z\leqslant 70$, $71\leqslant Z\leqslant 100$ and $101\leqslant Z\leqslant 120$, respectively. For an odd-$A$ nucleus, the blocking effect is considered with the equal filling approximation Li2012CPL ; Pan2022PRC . The above numerical conditions have been examined and determined in Refs. Pan2019IJMPE ; Zhang2020PRC ; Pan2022PRC .
2.2 Classical $r$-process
In this work, a site-independent $r$-process model, i.e., the so-called classical $r$-process model, is employed to specifically study the effect of nuclear mass on the $r$-process simulation. Classical $r$-process model can be regarded as a simplification of the dynamical $r$-process model, and has been successfully employed in describing $r$-process patterns of both the solar system and metal-poor stars Kratz1993APJ ; Kratz2007APJ ; Sun2008PRC ; Xu2013PRC . Nevertheless, it should be noted that the real neutron freeze-out after the equilibrium between neutron capture and photodisintegration reactions, as well as the fission recycling, are neglected in the present classical $r$-process model.
In the classical $r$-process model, iron group seed nuclei are irradiated by high-density neutron sources with a high temperature $T\gtrsim 1.5~\mathrm{GK}$. The $r$-process abundances are obtained by the superposition of abundances from the simulations in 16 different neutron flows with neutron densities in the range of $10^{20}-10^{27.5}\mathrm{~cm}^{-3}$. The weight $\omega$ and the irradiation time $\tau$ of each neutron flow follow exponential relations on neutron density $n_{n}$ Kratz1993APJ ; Chen1995PLB :